OpenStudy (specek18):

What is the simplified form of this expression? 8x3+24x2+18x−−−−−−−−−−−−−−√

OpenStudy (specek18):

here is the problem. will give medals, please help!! @AndrewTheCookie @ashbrunet @BaconBits @CaseyCarns @danish071996

OpenStudy (anonymous):

(x+1)(ax^2+bx+c)

OpenStudy (anonymous):

develope it

OpenStudy (specek18):

i'm confused on what to do for the first step...i'm sorry.

OpenStudy (anonymous):

the first one (choice)

OpenStudy (anonymous):

factor a \(2x\) out of the radicand, the result i think is a perfect square

OpenStudy (specek18):

okay i'm doing that now:)

OpenStudy (anonymous):

let me know what you got

OpenStudy (specek18):

wait factor the 2x out of all of the parts of the expression?

OpenStudy (anonymous):

factor a \(2x\) out of the radicand \[8x^3+24x^2+18x\]

OpenStudy (anonymous):

let me know what you get then we can do from there

OpenStudy (specek18):

4x^3+12x^2+9x is what i got...

OpenStudy (anonymous):

you divided by 2 maybe it is not clear what i mean by 'factor out the \(2x\)"

OpenStudy (anonymous):

\[8x^3+24x^2+18x=2x(something)\]

OpenStudy (specek18):

how do i factor the 2x out? i'm sorry, clearly having trouble...

OpenStudy (anonymous):

you do it mentally i think each term in \(8x^3+24x^2+18x\) has a factor of \(2x\) in it

OpenStudy (anonymous):

for example \(8x^3=2x\times 4x^2\) have you not factored out a common factor of a polynomial before?

OpenStudy (specek18):

i don't think i have....is this the same as foil?

OpenStudy (anonymous):

no foil is a stupid mnemonic device invented by dim math teachers to try to get you to remember the distributive law factoring is writing a sum as a product, not multiplying out

OpenStudy (anonymous):

this is how to factor, see if it rings a bell \[8x^3+24x^2+18x=2x(4x^2+12x+9)\]

OpenStudy (anonymous):

then you are supposed to recognize that \(4x^2+12x+9\) is a perfect square namely it is \[(2x+3)^2\]

OpenStudy (specek18):

okay...i see what you just did on the last post! about how to factor:)

OpenStudy (specek18):

that would make it the first option correct?

OpenStudy (anonymous):

so putting it all together eventually you get \[\sqrt{(2x+3)^2\times 2x}\]

OpenStudy (anonymous):

making the entire thing \[(2x+3)\sqrt{2x}\]

OpenStudy (specek18):

i am understanding slowly:) do you think you could help me with a few more problems?

OpenStudy (anonymous):

that would be the first option, but i am not so sure about the \(\pm\) part, but that is what you have

OpenStudy (anonymous):

sure i can help with another if i know it

OpenStudy (anonymous):

you want to post a screen shot?

OpenStudy (specek18):

i will do that, i'll also close this sothat i can give you another medal:)

OpenStudy (anonymous):

lol ok i don't really need another one though, but if you post a new question i will not have to scroll down so far

OpenStudy (specek18):

:)haha okay! either way:) and you deserve another.